# Accelerating Distributed Optimization: A Primal-Dual Perspective on Local Steps

### v1

@@ 3.1 A Primal-Dual Algorithm
This algorithm consists of two loops. In the inner loop, it performs multiple steps of Stochastic Gradient Descent on primal variables, each of which can be executed locally by each client. Subsequently, the coordinator broadcasts its variable, $x_{1}$, to all clients, who then update their dual variable, $\lambda_{m}$, by performing a single step of Gradient Ascent. All clients send their $\lambda_{m}$ back to the coordinator to start the next iteration.
@@ 3.2 Convergence Analysis for Centralized GA-MSGD
The theorem implies that GA-MSGD converges linearly, up to an error term due to the inexactness of the inner loops. Note that the Lagrangian is linear in $\lambda$, and the optimal complexity of solving a general minimax optimization problem that is only linear in the dual variable is $\mathcal{O}(\epsilon^{-\frac{1}{2}})$ [36]. Here linear convergence arises because we can show the dual function $\widetilde{\Psi}(\lambda)\triangleq\min_{x_{1},X}\widetilde{\mathcal{L}}(x_{1},X,\lambda)$ is strongly concave by leveraging the following proposition.
@@
Theorem 3 (Inner-loop Complexity).
@@
To find an $\epsilon$-optimal point $x$ such that $F(x)-F^{*}\leq\epsilon$, the total number of communication rounds in Algorithm 2 is
@@ 3.3 Catalyst Acceleration for GA-MSGD
For strongly convex objectives, it achieves the optimal communication complexity $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ [16, 51, 52] up to logarithmic terms for both deterministic ($\sigma=0$) and stochastic settings. The existing local method can only achieve such communication complexity asymptotically in the stochastic setting and needs to choose hyperparameters implicitly by solving an additional cubic equation [10]. Moreover, It improves the previous best-known gradient complexity of $\widetilde{\mathcal{O}}(\kappa^{\frac{3}{4}})$ for local methods [41] to near-optimal $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ in the deterministic setting, and attains a sample complexity nearly optimal in $\epsilon$ dependency for the stochastic setting. For convex objectives, it also achieves near-optimal $\widetilde{\mathcal{O}}(\epsilon^{-\frac{1}{2}})$ communication complexity [51]. To our knowledge, there are no existing local methods providing such complexities, even in the deterministic setting. For nonconvex objectives, it attains comparable complexity in the dependency of $\epsilon$ with state-of-the-art approaches such as [11, 4].
@@
The results presented above do not demonstrate linear speedup, as the term with variance $\sigma^{2}$ in the sample complexity does not decrease with the number of clients $M$. Existing local methods achieving linear speedup typically incorporate an interpolation step in the update of the primal variable [11, 4]. We defer the primal-dual interpretation of such a mechanism to future work. Additionally, there is potential for improvement in the dependency on $\kappa$ for the stochastic strongly convex setting and the dependency on $\epsilon$ for the stochastic convex setting. This improvement may hinge on designing better stopping criteria in subproblems for Catalyst. Recently, Lan and Li [17] designed a new stopping criterion using Stochastic Gradient Descent as a subroutine to achieve the optimal sample complexities, but it requires a tailored analysis for the subroutine. This approach could potentially be combined with our distributed algorithm in the future.
@@ 3.4 Catalyst Acceleration for SCAFFOLD
Combining SCAFFOLD with Catalyst not only realizes the advantages in Remark 1 but also has the $\sigma^{2}$ term in sample complexity diminishes linearly with the number of clients $M$.
@@ 4.1 A Primal-Dual Algorithm
Similar to Section 3.1, we implement Gradient Descent for $X$ in the inner loop, while utilizing Accelerated Gradient Ascent [34] for $\lambda$. This accelerated approach is employed to improve the dependency on $1-\sigma_{2}$.
@@
We present Acc-GA-MSGD in Algorithm 3, where $\nabla H$ is substituted with its stochastic estimate. An important observation is that the update of $X$ in the inner loops does not require communication.
@@
where $AX$ represents the consensus constraint for some matrix $A$ related to $W$. Examples of such algorithms include ADMM [25, 44], EXTRA [9, 45], and DIGing [32]. While this quadratic term enhances the convexity condition, it usually requires communication in the update of primal variables.
@@
Local Exact-Diffusion (LED), presented in 5, was recently introduced by Alghunaim [1]. It is equivalent to performing Gradient Ascent on the dual variable in the Lagrangian (4)—specifically, setting the momentum ($\beta$) to zero in Algorithm 3. In [1], LED achieves $\widetilde{\mathcal{O}}(\frac{\kappa^{2}}{1-\sigma_{2}})$ communication complexity for the strongly convex setting with sufficient local steps. We improve this to $\widetilde{\mathcal{O}}(\frac{\kappa}{1-\sigma_{2}})$ through our primal-dual interpretation of LED in Section C.4. According to Alghunaim [1], when the number of local steps is 1, LED reduces to Exact Diffusion [59], which is further equivalent to $\text{D}^{2}$ [48]. Therefore, we also provide a new primal-dual interpretation for these methods. By combining the idea from [7] with our analysis in the next subsection, we can recover the linear convergence for these methods in the strongly convex deterministic setting. However, this is not included in this paper because we focus on the local methods here.
@@
When the momentum ($\beta$) is set to zero, Algorithm 3 closely resembles decentralized Scaffnew [29], with the primary difference that Scaffnew employs a random number of inner-loop iterations. We also detail a centralized version of Algorithm 3 in Algorithm 6, which is almost equivalent to centralized Scaffnew, as discussed in Section C.5. Consequently, the control variate used in Scaffnew to mitigate client drift is interpreted as the dual variable in our Lagrangian formulation.
@@ 4.2 Convergence Analysis for Decentralized Acc-GA-MSGD
In contrast to our previous reformulation in the centralized setting, the coupling matrix between $X$ and $\lambda$ is no longer full-rank. However, we observe the following structure: $\lambda^{t}$ remains in $\operatorname{Span}(U)=\{U\hat{\lambda}:\hat{\lambda}\in\mathbb{R}^{M\times n}\}$ based on its update rule (5), as long as the initial $\lambda^{0}\in\text{Span}(U)$. Additionally, according to Lemma 3.1 in [45], an optimal $\lambda^{*}$ exists in Span($U$). We now introduce the key intermediate step of this work.
Theorem 5.
@@
This theorem can be viewed as a generalization of Proposition 1. In Lemma 3, we show that the dual function $\Psi(\lambda)\triangleq\min_{X}\mathcal{L}(X,\lambda)$ is $\frac{2M}{\mu}$-smooth and $\frac{M(1-\sigma_{2})}{L}$-strongly concave in $\text{Span}(U)$, which helps to establish of the following linear convergence.
@@
Theorem 7 (Inner-loop Complexity).
@@
Even with noisy gradients, Acc-GA-MSGD is able to achieve $\widetilde{\mathcal{O}}(\sqrt{\frac{\kappa}{1-\sigma_{2}}})$ communication complexity, which matches the lower bound for the deterministic setting  [42, 43]. Although Scaman et al. [42, 43] presented their lower bound for algorithms that perform communication after each step of gradient updates, their lower bound remains applicable as a communication complexity lower bound for a more general class of distributed zero-respecting algorithms that incorporate multiple local updates, as defined in Definition 1. This is because, with their worst-case function (see proof of Theorem 16 in [43] for more detail), performing multiple local steps does not activate more coordinates than a single local step.
Remark 5.
Our communication complexity has greatly improved upon the $\widetilde{\mathcal{O}}(\frac{\kappa^{2}}{1-\sigma_{2}})$ complexity for LED [1], and the $\widetilde{\mathcal{O}}(\frac{\kappa}{(1-\sigma_{2})c})$ complexity for stochastic Gradient Tracking with large enough minibatch [13, 39], where $c>\Theta(1-\sigma_{2})$ is another network-dependent constant.
Definition 1 (distributed zero-respecting algorithm [51]).
Let $\operatorname{supp}(v)=\left\{i\in[d]:v_{i}\neq 0\right\}$ and $\pi_{m}\left(t,m^{\prime}\right)$ be the last time before $t$ when machines $m$ and $m^{\prime}$ communicated. We say that an optimization algorithm is distributed zero-respecting if the $t$-th query on the $m$-th machine, $x_{m}^{t}$ satisfies
$\operatorname{supp}\left(x_{m}^{t}\right)\subseteq\bigcup_{s<t}\operatorname{supp}\left(\nabla F\left(x_{m}^{s}\right)\right)\cup\bigcup_{m^{\prime}\neq m}\bigcup_{s\leq\pi_{m}\left(t,m^{\prime}\right)}\operatorname{supp}\left(\nabla F\left(x^{s}_{m^{\prime}}\right)\right).$
@@ 4.3 Acceleration through Catalyst
  • (NC) comm: $\widetilde{\mathcal{O}}\left(\frac{1}{\sqrt{1-\sigma_{2}}\epsilon^{2}}\right)$;  sample: $\widetilde{\mathcal{O}}\left(\frac{1}{\epsilon^{2}}+\frac{\sigma^{2}}{(1-\sigma_{2})^{3}\epsilon^{4}}\right)$.
@@
For strongly convex objectives, this further improves the dependency on $\kappa$ in sample complexity compared with Acc-GA-MSGD. For stochastic convex objectives, we achieve a near-optimal communication complexity of $\widetilde{\mathcal{O}}((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-\frac{1}{2}})$. As noted in Remark 4, the lower bound of $\Omega((1-\sigma_{2})^{-\frac{1}{2}}\epsilon^{-\frac{1}{2}})$ in [43] applies as communication lower bound for distributed zero-respecting algorithms. Our communication complexity improves over the existing state-of-the-art complexity of $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-1})$  [1]. For stochastic nonconvex objectives, our method also shows improvement in communication complexity over the best-known results: $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-2})$ from [1] and $\mathcal{O}((1-\sigma_{2})^{-2}\epsilon^{-2})$ from [27], particularly in terms of the dependence on $1-\sigma_{2}$. It is important to note that these comparisons of communication complexities assume that a sufficient number of local steps are conducted in each round. However, scenarios where only a few local steps are permitted are beyond the scope of this paper.
@@ 5 Conclusion and Future Directions
This work demonstrates that communication complexities comparable to those in deterministic settings can be achieved in stochastic settings using local methods for both centralized and decentralized problems. Additionally, we establish a link between the Lagrangian formulation of distributed optimization and minimax optimization. This connection paves the way for applying extensive minimax optimization literature, including Gradient Descent Ascent (GDA)-type algorithms [24, 56, 57, 3, 60, 55], to distributed optimization.
@@ C.1 Useful Lemmas
Proof for Proposition 1 and Theorem 5
@@ C.2 Proofs for Acc-GA-MSGD
    Under Assumption 3 for $W$, with $U=\sqrt{I-W}$, we know that the largest singular value of $U$ is no larger than $\sqrt{2}$, and the smallest nonzero singular value of $U$ is no smaller than $\sqrt{1-\sigma_{2}}$. We conclude by applying Proposition 1 and Theorem 5. ∎
@@ C.4 Better Communication Complexity for LED
    where in the inequality we use Proposition 5 and a equivalent definition of strong concavity. With $\tau_{2}=\frac{\mu}{2M}$,

### v2

@@ 3.1 A Primal-Dual Algorithm
This algorithm consists of two loops. In the inner loop, it performs multiple steps of Stochastic Gradient Descent (SGD) on primal variables, each of which can be executed locally by each client. Subsequently, the coordinator broadcasts its variable, $x_{1}$, to all clients, who then update their dual variable, $\lambda_{m}$, by performing a single step of Gradient Ascent. All clients send their $\lambda_{m}$ back to the coordinator to start the next iteration.
While this algorithm is described using Stochastic Gradient Descent (SGD) to solve for the primal variables, formula (3) allows for considerable flexibility in selecting optimization methods. Alternatives to SGD, such as adaptive methods like Adam [20] or Shampoo [15], can be employed. Furthermore, it is possible for each client to use different optimizers in local steps.
@@ 3.2 Convergence Analysis for Centralized GA-MSGD
The theorem implies that GA-MSGD converges linearly, up to an error term due to the inexactness of the inner loops. Note that the Lagrangian is linear in $\lambda$, and the optimal complexity of solving a general minimax optimization problem that is only linear in the dual variable is $\mathcal{O}(\epsilon^{-\frac{1}{2}})$ [49]. Here linear convergence arises because we can show the dual function $\widetilde{\Psi}(\lambda)\triangleq\min_{x_{1},X}\widetilde{\mathcal{L}}(x_{1},X,\lambda)$ is strongly concave by leveraging the following proposition. A similar argument is presented in [11, 13].
@@
Theorem 2 shows that the algorithm will still converge even if optimizers other than SGD are used in the inner loop, as long as the required accuracy for these loops is achieved. For simplicity, we present the inner-loop complexity when every client uses SGD as local steps, following the SGD convergence for strongly convex function [59].
Lemma 1 (Inner-loop Complexity).
@@
To find an $\epsilon$-optimal point $x$ such that $\mathbb{E}\|F(x)-F^{*}\|\leq\epsilon$, the total number of communication rounds in Algorithm 2 is
@@ 3.3 Catalyst Acceleration for GA-MSGD
For strongly convex objectives, it achieves the optimal communication complexity $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ [64, 65] up to logarithmic terms for both deterministic ($\sigma=0$) and stochastic settings. The existing local method can only achieve such communication complexity asymptotically in the stochastic setting [18]. Moreover, It improves the previous best-known gradient complexity of $\widetilde{\mathcal{O}}(\kappa^{\frac{3}{4}})$ for local methods [54] to near-optimal $\widetilde{\mathcal{O}}(\sqrt{\kappa})$ in the deterministic setting, and attains a sample complexity nearly optimal in $\epsilon$ for the stochastic setting. For convex objectives, it also achieves near-optimal $\widetilde{\mathcal{O}}(\epsilon^{-\frac{1}{2}})$ communication complexity [64]. To our knowledge, there are no existing local methods providing such complexities. For nonconvex objectives, it attains comparable complexity in the dependency of $\epsilon$ with state-of-the-art approaches such as [19, 8].
@@
The results above do not demonstrate linear speedup, as the term with variance $\sigma^{2}$ in the sample complexity does not decrease with the number of clients $M$. Existing local methods achieving linear speedup typically incorporate an interpolation step in the update of the primal variable [19, 8]. Additionally, there is room for improvement in sample complexities in the dependency on $\kappa$ for the strongly convex setting and the dependency on $\epsilon$ for the convex setting. This improvement may hinge on designing better stopping criteria in subproblems for Catalyst. Recently, Lan and Li [26] designed a new stopping criterion using SGD as a subroutine to achieve the optimal sample complexities, but it requires a tailored analysis for the subroutine. The approach could potentially be combined with our distributed algorithm in the future.
@@ 3.4 Catalyst Acceleration for SCAFFOLD
Combining SCAFFOLD with Catalyst not only realizes the advantages in Remark 1 but also has the $\sigma^{2}$ term in sample complexity diminishes linearly with the number of clients $M$. However, compared to GA-MSGD, SCAFFOLD requires smaller step sizes for local updates (see Section C in Woodworth et al. [64]) and offers less flexibility in choosing the local optimizer.
@@ 4.1 A Primal-Dual Algorithm
This formula is widely used in the literature [28, 27, 69]. Similar to Section 3.1, we implement Gradient Descent for $X$ in the inner loop, while utilizing Accelerated Gradient Ascent [47] for $\lambda$. This accelerated approach is employed to improve the dependency on $1-\sigma_{2}$.
@@
We present Acc-GA-MSGD in Algorithm 3, where $\nabla H$ is substituted with its stochastic estimate. An important observation is that the update of $X$ in the inner loops does not require communication. Substituting $X^{t,K}$ with $X^{*}(\zeta^{t})\triangleq\operatorname*{arg\,min}_{X}H(X)+\langle\zeta^{t},X\rangle$ in the update of $\zeta^{t+1}$ transforms the algorithm into Accelerated Dual Ascent [55]. When the dual gradient $X^{*}(\lambda)$ is not directly accessible, Uribe et al. [62] have employed Accelerated Gradient Descent in the inner loops to approximate it. Following a similar concept, Algorithm 3 employs Stochastic Gradient Descent (SGD) in the inner loops to achieve this approximation. Since the update of $X$ is decomposable across clients, each client may also independently select their optimization method, such as ADAM [20].
@@
where $AX$ represents the consensus constraint for some matrix $A$ related to $W$. Examples of such algorithms include EXTRA [17, 57] and DIGing [45]. While this quadratic term enhances the convexity condition, it usually requires communication in the update of primal variables.
@@
Local Exact-Diffusion (LED), presented in 5, was recently introduced by Alghunaim [1]. It is equivalent to performing Gradient Ascent on the dual variable in the Lagrangian (4)—specifically, setting the momentum ($\beta$) to zero in Algorithm 3. In [1], LED achieves $\widetilde{\mathcal{O}}(\frac{\kappa^{2}}{1-\sigma_{2}})$ communication complexity for the strongly convex setting with sufficient local steps. We improve this to $\widetilde{\mathcal{O}}(\frac{\kappa}{1-\sigma_{2}})$ through the primal-dual interpretation of LED in Section C.4. According to Alghunaim [1], when the number of local steps is 1, LED reduces to Exact Diffusion [74] and $\text{D}^{2}$ [60].
@@
When the momentum ($\beta$) is set to zero, Algorithm 3 closely resembles decentralized Scaffnew [40], with the primary difference that Scaffnew employs a random number of inner-loop iterations. We also detail a centralized version of Algorithm 3 in Algorithm 6, which is almost equivalent to centralized Scaffnew, as discussed in Section C.5. Consequently, the control variate used in Scaffnew to mitigate client drift is interpreted as the dual variable in the Lagrangian formulation.
@@ 4.2 Convergence Analysis for Decentralized Acc-GA-MSGD
In contrast to our previous reformulation in the centralized setting, the coupling matrix between $X$ and $\lambda$ is no longer full-rank. However, we observe the following structure: $\lambda^{t}$ remains in $\operatorname{Span}(U)=\{U\hat{\lambda}:\hat{\lambda}\in\mathbb{R}^{M\times n}\}$ based on its update rule (5), as long as the initial $\lambda^{0}\in\text{Span}(U)$. Additionally, according to Lemma 3.1 in [57], an optimal $\lambda^{*}$ exists in Span($U$). In the following, we present an intermediate step.
Lemma 2.
@@
A similar step is also described in [55], but it assumes the existence of the Hessian of the Fenchel conjugate of $f$. This can be met when $f$ is twice differentiable. The lemma can be viewed as a generalization of Proposition 1. In Lemma 6, we show that the dual function $\Psi(\lambda)\triangleq\min_{X}\mathcal{L}(X,\lambda)$ is $\frac{2M}{\mu}$-smooth and $\frac{M(1-\sigma_{2})}{L}$-strongly concave in $\text{Span}(U)$, which helps to establish of the following linear convergence.
@@
Lemma 3 (Inner-loop Complexity).
@@
Acc-GA-MSGD achieves $\widetilde{\mathcal{O}}\left(\sqrt{\frac{\kappa}{1-\sigma_{2}}}\right)$ communication complexity, which matches the lower bound for the deterministic setting  [55, 56]. Compared with other stochastic algorithms, our communication complexity has improved upon the $\widetilde{\mathcal{O}}\left(\frac{\kappa^{2}}{1-\sigma_{2}}\right)$ complexity for LED [1], and the $\widetilde{\mathcal{O}}\left(\frac{\kappa}{(1-\sigma_{2})c}\right)$ complexity for stochastic Gradient Tracking with large enough minibatch [22, 52], where $c>\Theta(1-\sigma_{2})$ is another network-dependent constant.
@@ 4.3 Acceleration through Catalyst
  • (NC) comm: $\widetilde{\mathcal{O}}\left(\frac{1}{\sqrt{1-\sigma_{2}}\epsilon^{2}}\right)$;  sample: $\widetilde{\mathcal{O}}\left(\frac{1}{\sqrt{1-\sigma_{2}}\epsilon^{2}}+\frac{\sigma^{2}}{(1-\sigma_{2})^{3}\epsilon^{4}}\right)$.
@@
The communication complexities in strongly convex, convex, and nonconvex settings match with the lower bounds in [55, 56, 37], respectively, up to logarithmic terms. Although the sample complexities include the dependency in $1-\sigma_{2}$, it can be eliminated by Chebyshev acceleration, detailed in Section 4.2 of [55] and widely adopted in the literature [69, 25]. In the stochastic convex setting, our communication complexity improves over the $\mathcal{O}((1-\sigma_{2})^{-1}\epsilon^{-1})$ complexity in [1]. In the stochastic nonconvex setting, the existing near-optimal algorithm DeTAG [37] requires large minibatches.
@@ 5 Conclusion and Future Directions
This work demonstrates that communication complexities comparable to those in deterministic settings can be achieved by stochastic local methods without using minibatches of large size. The connection between the primal-dual algorithms and local steps paves the way for applying extensive minimax optimization literature, including Gradient Descent Ascent (GDA)-type algorithms [33, 71, 72, 5, 75, 70], to distributed optimization. Furthermore, it is intriguing to investigate the potential benefits of allowing each client to adopt adaptive optimization algorithms for distributed optimization within the primal-dual framework.
@@ C.1 Useful Lemmas
Proof for Proposition 1 and Lemma 2
@@ C.2 Proofs for Acc-GA-MSGD
    Under Assumption 3 for $W$, with $U=\sqrt{I-W}$, we know that the largest singular value of $U$ is no larger than $\sqrt{2}$, and the smallest nonzero singular value of $U$ is no smaller than $\sqrt{1-\sigma_{2}}$. We conclude by applying Proposition 1 and Lemma 2. ∎
@@ C.4 Better Communication Complexity for LED
    where in the inequality we use Lemma 2 and a equivalent definition of strong concavity. With $\tau_{2}=\frac{\mu}{2M}$,